Consider two streams of cash flows, A and B. Stream A’s first cash flow is $9,600 and is received three years from today. Future cash flows in Stream A grow by 3 percent in perpetuity. Stream B’s first cash flow is −$9,300, is received two years from today, and will continue in perpetuity. Assume that the appropriate discount rate is 11 percent.

  

a.

What is the present value of each stream? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

  

  Present value
  Stream A $   
  Stream B $   

  

b.

Suppose that the two streams are combined into one project, called C. What is the IRR of Project C? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  

  IRR %

    

c.

What is the correct IRR rule for Project C?

Accept the project if the discount rate is above the IRR.
Accept the project if the discount rate is below the IRR.
Accept the project if the discount rate is equal the IRR.
Answer #1

1.
=9600/(11%-3%)*1/1.11^2=97394.6919892866

2.
=-9300/11%*1/1.11=-76167.0761670762

3.
-9300/IRR*(1-1/(1+IRR))+9600/(IRR-3%)*1/(1+IRR)^2=0

=>IRR=-164.64%, 64.638%

4.
Accept the project if the discount rate is below the IRR.